OpenStudy (anonymous):

826.45 moles of Na produce how many moles of Ti?

OpenStudy (anonymous):

that's problem from MIT course right? 4Na + TiCl4 -> Ti + 4NaCl n(Ti) / n(Na) = 1/4 n(Ti) = 1/4 * n(Na)